The product of two consecutive positive even numbers is 16 greater than 16 times of the odd number between them. Find the two even numbers.
Answer by CubeyThePenguin(3113)

consecutive positive even integers: x, (x+2)

x(x+2) = 16 + 16(x+1)
x^2 + 2x = 16x + 32
x^2 - 14x - 32 = 0
(x - 16)(x + 2) = 0
x = 16, x = -2

x is positive, so the integers are 14, 16, and 18.